# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v3

@@ 1 Introduction
To address these issues, we propose MixGRPO, a method that optimizes a strategic subset of denoising steps to drastically reduce computational overhead while ensuring a more focused and efficient optimization. Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest.
@@ 3.1 Mixed ODE-SDE Sampling in GRPO
where the SDE drift term for the sampling process is defined as $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\triangleq\mathbf{v}_{\theta}(\mathbf{x}_{t},t)+\frac{{\sigma^{2}_{t}}\left(\mathbf{x}_{t}+\left(1-t\right)\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\right)}{2t}$. According to Eq. (1), we only need to optimize GRPO [32] at the time within the interval $S$ for $N$ samples in the group. For an unbiased advantage estimation in GRPO, we constrain the source of variation across images generated from a single prompt. Specifically, we fix the position of the SDE window, thereby sourcing all stochasticity from an identical timestep range in a GRPO group. This not only concentrates the gradients within this window but is also essential for preserving the unbiased nature of the optimization. The final training objective is represented as follows:
@@ 4.2 Implementation Details
For training, all experiments are conducted on 32 Nvidia GPUs with a batch size per gpu of 1 and a maximum of 300 iterations. We use AdamW [20] as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at fp32 precision.
@@ 11 Cross-Dataset Experiments
The comparative results for the window size ($w$) are presented in Table 13. It is evident that both $w=4$ and $w=6$ represent optimal settings. Specifically, $w=6$ slightly surpasses $w=4$ only on the in-domain test set of Cross-dataset Experiment 1, and in other experiments, it only holds a marginal advantage in the HPS-v2.1 [40] score. It is crucial to note that the window size $w$ corresponds to the number of denoising timesteps that must be optimized during each training iteration, which, in turn, is linearly correlated with the overall optimization overhead of the Reinforcement Learning (RL) process. Consequently, $w=4$ is identified as the best compromise, offering an optimal trade-off between performance and computational efficiency. Regarding parameter sensitivity, we observe that the change in performance metrics—across both in-domain and out-of-domain tests—is relatively minimal when $w$ varies between $2$ and $6$. This stability suggests that the model exhibits robust performance with respect to the window size setting. As long as the window size remains within a reasonable range (e.g., $2\leq w\leq 6$), the model is capable of effective learning and maintaining stable performance, thereby reducing necessity for meticulous hyperparameter tuning.

### v4

@@ 1 Introduction
To address these issues, we propose MixGRPO, a method that optimizes a strategic subset of denoising steps to drastically reduce computational overhead while ensuring a more focused and efficient optimization. Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest. Recently, Coefficients-Preserving Sampling (CPS) [37] was proposed as a more principled alternative to standard SDE sampling. Further details are provided in 12.
@@ 3.1 Mixed ODE-SDE Sampling in GRPO
where the SDE drift term for the sampling process is defined as $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\triangleq\mathbf{v}_{\theta}(\mathbf{x}_{t},t)+\frac{{\sigma^{2}_{t}}\left(\mathbf{x}_{t}+\left(1-t\right)\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\right)}{2t}$. According to Eq. (1), we only need to optimize GRPO [32] at the time within the interval $S$ for NN samples in the group. For an unbiased advantage estimation in GRPO, we constrain the source of variation across images generated from a single prompt. Specifically, we fix the position of the SDE window, thereby sourcing all stochasticity from an identical timestep range in a GRPO group. This not only concentrates the gradients within this window but is also essential for preserving the unbiased nature of the optimization. The final training objective is represented as follows:
@@ 4.2 Implementation Details
For training, all experiments are conducted on 32 Nvidia GPUs with a batch size per gpu of 1 and a maximum of 300 iterations. We use AdamW [20] as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at fp32 precision. We also tried CPS sampling instead of normal SDE sampling, the details can seen in 12 .
@@ 11 Cross-Dataset Experiments
The comparative results for the window size ($w$) are presented in Table 13. It is evident that both $w=4$ and $w=6$ represent optimal settings. Specifically, $w=6$ slightly surpasses $w=4$ only on the in-domain test set of Cross-dataset Experiment 1, and in other experiments, it only holds a marginal advantage in the HPS-v2.1 [41] score. It is crucial to note that the window size $w$ corresponds to the number of denoising timesteps that must be optimized during each training iteration, which, in turn, is linearly correlated with the overall optimization overhead of the Reinforcement Learning (RL) process. Consequently, $w=4$ is identified as the best compromise, offering an optimal trade-off between performance and computational efficiency. Regarding parameter sensitivity, we observe that the change in performance metrics—across both in-domain and out-of-domain tests—is relatively minimal when $w$ varies between 22 and 66. This stability suggests that the model exhibits robust performance with respect to the window size setting. As long as the window size remains within a reasonable range (e.g., $2\leq w\leq 6$), the model is capable of effective learning and maintaining stable performance, thereby reducing necessity for meticulous hyperparameter tuning.
@@
12 Coefficients-Preserving Sampling
In our MixGRPO framework, introducing stochasticity during the inference phase is crucial for effective exploration in reinforcement learning. While a common practice involves the use of Stochastic Differential Equations (SDEs), we adopt Coefficients-Preserving Sampling (CPS) [37] as a more refined alternative to maintain the integrity of the probability path.
The standard SDE-based sampling, often discretized through the Euler-Maruyama method, follows the update rule:
$\mathbf{x}_{t_{i-1}}=\mathbf{x}_{t_{i}}-\mathbf{v}_{t_{i}}(\mathbf{x}_{t_{i}},t_{i})\Delta t+\sqrt{2\sigma^{2}\Delta t}\mathbf{\epsilon}_{i},\qquad\mathbf{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (24)
Although this formulation provides the necessary stochasticity, it tends to inject excessive independent noise at each discretization step, resulting in ”grainy” artifacts in the generated samples. As observed in our preliminary experiments, such high-frequency noise can lead to reward hacking, where reward models (e.g., Pick Score [12] or HPS-v2.1 [41]) prioritize low-level textures over structural coherence, thereby hindering the convergence of RL optimization.
To address these issues, we utilize CPS [37], which reformulates the transition process by drawing inspiration from DDIM [33] framework. The update rule for CPS [37] is defined as:
$\mathbf{x}_{t_{i-1}}=\frac{1-t_{i-1}}{1-t_{i}}\mathbf{x}_{t_{i}}+\left(t_{i-1}-\frac{1-t_{i-1}}{1-t_{i}}t_{i}\right)\mathbf{v}_{t_{i}}+\sigma_{t_{i}}\mathbf{\epsilon}_{i}$ (25)
The core advantage of CPS [37] lies in its strategic construction of coefficients for $\mathbf{x}_{t_{i}}$ and $\mathbf{v}_{t_{i}}$, which preserves the linear interpolation structure of Flow Matching while introducing controlled stochasticity via $\sigma_{t_{i}}$. By effectively eliminating sampling artifacts inherent in SDE-based methods, CPS [37] yields cleaner images that provide more reliable feedback for MixGRPO. In our implementation, we set $NFE_{\pi_{\theta_{old}}}=25$, window size $w=4$, and window stride $s=1$ with fixed initial noise, employing HPS-v2.1 [41], Pick Score [12], and ImageReward [42] as multi-reward metrics. Both the original SDE sampling and CPS [37] sampling were trained for 300 steps and evaluated on the HPDv2 dataset [41]. As demonstrated in Table 15, the integration of CPS [37] leads to superior performance across all reward metrics, with qualitative comparisons in Figure 9 further illustrating its enhanced visual fidelity.
Table 15: Comparison with CPS sampling
Method Pick Score ImageReward HPS-v2.1
FLUX 0.2265 1.0883 0.3126
MixGRPO-SDE 0.2344 1.5901 0.3648
MixGRPO-CPS 0.2382 1.6449 0.3692
@@ 13 More Visualized Results
Figure 9: Comparison of the visualization results of FLUX, SDE sampling and CPS sampling